A plane leaves an airport at 3 P.M. At 4 P.M., another plane leaves the same airport traveling in the same direction at a speed that is 150 mph faster than that of the first plane. Four hours after the first plane takes off, the second plane is 250 mi ahead of the first plane. How far did the second plane travel?
Answer by mananth(16949)

A plane leaves an airport at 3 P.M. At 4 P.M., another plane leaves the same airport traveling in the same direction at a speed that is 150 mph faster than that of the first plane. Four hours after the first plane takes off, the second plane is 250 mi ahead of the first plane. How far did the second plane travel?
First plane speed = x
second plane speed = x+150
Distance traveled by second plane - distance by first plane = 250 miles
distance traveled by first plane in 4 hours = 4x
distance traveled by second plane = 3(x+150) ( 1 hour later)
3(x+150)-4x=250
3x+450-4x=250
-x=250-450
-x=-200
x= 200
speed of first plane = 200 mph
speed of second plane = 200+150 = 350 mph
distance of second plane = 3*350=1050 miles